A defined-benefit retirement plan – once you retire - will pay $30,000 per year for 40 years and increase the annual payment by three-percent each year. What is the present value at retirement if the discount rate is 7 percent?

$265,311

$381,212

$456,713

$586,620

$642,311

Please help urgently. No work needs to be shown.
Answer #1

First payment $30,000 present value = p/(r-g) x 1-[(1+g)/(1+r)]n p/(r-g) = $750,000 30000/(0.07-0.03) 1-[(1+g)/(1+r)]n = 0.78